# thanks: co-first authorsthanks: co-first authors

### v1

@@ Abstract
Recent advances in neural recording technology allow simultaneously recording action potentials from hundreds to thousands of neurons in awake, behaving animals. However, characterizing spike patterns in the resulting data, and linking these patterns to behaviour, remains a challenging task. The lack of a rigorous mathematical language for variable numbers of events (spikes) emitted by multiple agents (neurons) is an important limiting factor. We introduce a new mathematical operation to decompose complex spike patterns into a set of simple, structured elements. This creates a mathematical language that allows comparing spike patterns across trials, detecting sub-patterns, and making links to behaviour via a clear distance measure. We apply the method to dual Utah array recordings from macaque prefrontal cortex, where this technique reveals previously unseen structure that can predict both memory-guided decisions and errors in a virtual-reality working memory task. These results demonstrate that this technique provides a powerful new approach to understand structure in the spike times of neural populations, at a scale that will continue to grow more and more rapidly in upcoming years.
@@
By solving the key technical problem with the inner product, the mathematical framework we introduce here provides a novel technique to analyze the large-scale spike recordings that are becoming increasingly important in neuroscience. We demonstrate the utility of this approach by providing example applications to large-scale recordings from macaque monkeys. Notably, applying this method to spikes from prefrontal cortex (PFC) allows predicting both choices and errors during a virtual reality (VR) working memory task. Taken together, these initial applications to large-scale spike train recordings demonstrate that this technique provides a powerful new approach to understand structure in the spike times of neural populations, at a scale that will continue to grow more and more rapidly in upcoming years.
@@ I.2 Multi-sample Discrete Helix Transform (ms-DHT)
In summary, the ms-DHT takes a pattern of spikes across many neurons (Fig. 3a), and breaks it down into a set of structured sub-patterns (Fig. 3b), called helices, defined in terms of an underlying cylindrical geometry (Fig. 3c). We term the result of this decomposition for a single helix (Fig. 3b, red dot) the contribution of that helix to the original spike pattern because it measures the strength and timing of the match between that helix and the original pattern. The vector of contributions across the full set of helices corresponds to the multi-sample Discrete Helix Transform (ms-DHT). This vector (depicted in the complex plane in Fig. 3) provides a description of the fine-scale structure in a spike pattern, which we term its fingerprint. Note that fingerprints are unique for a given labeling of the neurons (see Methods Sec. II.10); however, relationships between fingerprints are preserved if we re-label the neurons. This means that an identified link to behavior is not specific to a specific labeling, and that the link will not be impacted by a re-ordering of the cells. A key feature of the ms-DHT, due to the generalized inner product, is that the decomposed spike sequences can overlap in complex patterns across both neurons and time (Fig. 3d).
I.3 Spike patterns in macaque prefrontal cortex
@@
The recordings from two Utah arrays yield between 200-300 well-isolated single units in each session. In previous work, we found that sequences of single neuron activations encoded working memory content in this task (Fig. 4b) [37]. These sequences were composed of brief 100-200 ms elevations of firing rate well above baseline, and the position of each cell in the sequence was determined by the time it reached its maximum firing rate. The sequences had similar ordering and timing on trials of the same condition, and were separable between conditions. While we previously restricted to consider only one spiking event per cell on each trial, the ms-DHT now allows us to consider the complex pattern of all such elevations across all cells within a trial (Fig. 4c, Methods Sec.  II.9). This allows us to determine (1) whether we can replicate our previous results with this method and (2) whether including additional spiking events provides any additional insight.
@@
To what do these “misplaced” trials correspond? When studying the navigation paths subjects took on these “misplaced” trials, the navigation paths are at first more similar to the trials with which the spike pattern clustered. Taking for example an individual trial with a target in the left column (encoded in blue), the spike pattern during navigation was more similar to trials with a target in the right column (Fig. 4g, blue dot in yellow cluster, circled in black). On that trial, the subject initially navigated to the right (Fig. 4h, black path compared to blue and yellow paths), reached the end of the arena without receiving a reward, then turned around and headed to the left to correct the mistake, eventually reaching the correct target (bold circle). These results extend to incorrect trials, where the similarity between these incorrect trials and the spike patterns on correct trials could explain the specific type of error made on individual trials (Fig. 4i,j). Finally, these results were consistent across trials, where we found that paths from misclassified trials were significantly more similar to paths in the predicted cluster than to paths in their true cluster (STATS) (see Supplement Fig. S3). These results show that spike patterns occurring during navigation could more closely reflect the navigation strategy than the end of the trajectory at the correct target (Fig. 4h,j, bold circle), and that analysis of spike patterns by the ms-DHT can predict the specific errors made on individual trials (see Methods Sec. II.9 for details).
@@ Discussion
In this work, we have introduced a novel framework for characterizing patterns of spiking activity across large populations of neurons. This approach represents the first complete mathematical language for spike patterns, which provides a simple way to describe, compare, and decompose patterns of spikes on a variety of timescales. We provide example applications to large-scale spike recordings from macaque monkeys. These applications demonstrate how the multi-sample Discrete Helix Transform (ms-DHT) can be used to find behaviourally relevant structure in spike patterns that can then be directly related to the spikes on a single trial level.
@@ II.6 Dimensionality reduction procedure
In more detail, to compare $N_{T}$ spike patterns across $N$ cells, we begin by computing $\Psi_{H}(s_{i})$ for the spike pattern $s$ on each trial $i\in[1,N_{T}]$. We next compute the distance between each pair of spike patterns to produce the distance matrix $D\in\mathcal{M}_{N_{T}\times N_{T}}$ with $d_{i,j}:=||\Psi_{H}(s_{i})-\Psi_{H}(s_{j})||$. We then compute the eigenvalues $\lambda_{1},...\lambda_{N_{T}}$ and corresponding eigenvectors $\vec{v}_{1},....\vec{v}_{N_{T}}$, with eigenvalues ordered such that $|\lambda_{1}|\geq|\lambda_{2}|\geq...\geq\lambda_{N_{T}}$. Finally, we obtain the the $n$-dimensional projection $P:=D[\vec{v}_{1},...\vec{v}_{n}]$. This projection underlies the clustering in Fig. 4d,g,i.
@@ II.8 Significantly different helices
To determine whether helix contributions differed across trial conditions (as in Fig. 4n), we first computed a set of $1000$ bootstrapped centroids for each condition. The helix was considered to differ between conditions if these sets of centroids were non-overlapping (see for example Supplement Fig. S2b).
II.9 Application to Utah array data
@@
Experimental details and information about neuronal activation sequences can be found in [37]. To extend these sequences to more complicated spike patterns, spike times were defined to be the center of a spiking event when a neuron exceeded $80\%$ of its peak firing rate. Epochs were divided by task timing: ITI = $1$ second before the appearance of the cue; “cue” = $3$ seconds following cue onset, while the visual cue is present on screen; “delay” = $2$ second memory delay; “response” = first $2$ seconds of navigation.
@@
We hypothesized that the paths taken by the animal on incorrectly clustered trials would more closely resemble the paths in the predicted cluster as opposed to the true cluster (i.e., if a blue trial is grouped into the yellow cluster, we expect the path on that trial to look more like a yellow path “predicted cluster” than a blue path “true cluster” - see Supplement Figs. S3a,b). We used the Fréchet distance to measure the distance between paths. As a baseline, we computed the distance between paths on trials correctly assigned to the same true cluster (see Supplement Fig. S3c “correct to true cluster”). This provides a value for the distance between the paths on all correct yellow column trials that fall into the yellow column cluster, for example. For each incorrectly clustered trial, we then computed the mean similarity to other paths in its predicted cluster, compared to other paths in its true cluster. As expected, paths on incorrectly clustered trials were significantly more similar to the paths on trials in the predicted cluster than to other trials of the same condition (true cluster) (p¡0.003, see Supplement Fig. S3). Together, this shows that transient spike patterns in the navigation epoch of the task closely reflect the navigation strategy taken by the subject on individual trials – so closely, in fact, that the observed spike pattern can predict the specific type of error made on individual trials.
II.10 Mathematical Details
II.10.1 Definitions

### v2

@@ Abstract
Recent advances in neural recording technology allow simultaneously recording action potentials from hundreds to thousands of neurons in awake, behaving animals. However, characterizing spike patterns in the resulting data, and linking these patterns to behaviour, remains a challenging task. The lack of a rigorous mathematical language for variable numbers of events (spikes) emitted by multiple agents (neurons) is an important limiting factor. We introduce a new mathematical operation to decompose complex spike patterns into a set of simple, structured elements. This creates a mathematical language that allows comparing spike patterns across trials, detecting sub-patterns, and making links to behaviour via a clear distance measure. We first demonstrate the method using Neuropixel recordings from macaque motor cortex. We then apply the method to dual Utah array recordings from macaque prefrontal cortex, where this technique reveals previously unseen structure that can predict both memory-guided decisions and errors in a virtual-reality working memory task. These results demonstrate that this technique provides a powerful new approach to understand structure in the spike times of neural populations, at a scale that will continue to grow more and more rapidly in upcoming years.
@@
By solving the key technical problem with the inner product, the mathematical framework we introduce here provides a novel technique to analyze the large-scale spike recordings that are becoming increasingly important in neuroscience. We demonstrate the utility of this approach by providing example applications to Neuropixel and Utah array data from macaque monkeys. Notably, applying this method to spikes from prefrontal cortex (PFC) allows predicting both choices and errors during a virtual reality (VR) working memory task. Taken together, these initial applications to large-scale spike train recordings demonstrate that this technique provides a powerful new approach to understand structure in the spike times of neural populations, at a scale that will continue to grow more and more rapidly in upcoming years.
@@ I.2 Multi-sample Discrete Helix Transform (ms-DHT)
In summary, the ms-DHT takes a pattern of spikes across many neurons (Fig. 3a), and breaks it down into a set of structured sub-patterns (Fig. 3b), called helices, defined in terms of an underlying cylindrical geometry (Fig. 3c). We term the result of this decomposition for a single helix (Fig. 3b, red dot) the contribution of that helix to the original spike pattern because it measures the strength and timing of the match between that helix and the original pattern. The vector of contributions across the full set of helices corresponds to the multi-sample Discrete Helix Transform (ms-DHT). This vector (depicted in the complex plane in Fig. 3) provides a description of the fine-scale structure in a spike pattern, which we term its fingerprint. Note that fingerprints are unique for a given labeling of the neurons (see Methods Sec. II.11); however, relationships between fingerprints are preserved if we re-label the neurons. This means that an identified link to behavior is not specific to a specific labeling, and that the link will not be impacted by a re-ordering of the cells. A key feature of the ms-DHT, due to the generalized inner product, is that the decomposed spike sequences can overlap in complex patterns across both neurons and time (Fig. 3d).
I.3 Spike patterns in macaque motor cortex
As a first demonstration of applying this method to experimental data, we consider Neuropixel recordings from macaque area M1 during a motor task [36]. The task set up is depicted in Fig. 4a. On each trial, a robotic exoskeleton applies a perturbation to the subject’s arm in one of two directions, chosen at random with equal probability. These perturbations cause either an extension or a flexion of the elbow (Fig.  4b), and occur at a random time following the start of the trial. In the 100ms following these large, salient perturbations, the average firing rate across neurons in this area increases transiently (Fig. 4c). Activity in area M1 during this short timescale response is known to encode this type of salient perturbation [37], and during this time, the subject instinctively corrects for the elbow load to return their arm to the starting location. In what follows, we decompose and analyze the spike patterns during this $100$ms transient response to the perturbations. This represents a clear and simple demonstration of the approach in a dataset where differences in spiking activity are expected to occur across conditions [37].
[…]
On each trial, we study the set of spikes immediately following the perturbation (Fig. 4d, blue to dashed black line). We take the ms-DHT of the spike pattern, and study the resulting helix contribution. We found that some helices were detected more frequently on trials of a specific perturbation direction (i.e. $||\mu_{k}||$, for a specific $k$, was high relative to chance level for one condition). For example, in this session, helix 2 (Fig. 4d-h) was detected more often following extension perturbations (blue), while helix 8 (Fig. 4i,j) was detected more often on flexion trials (pink). To visualize this structure, we can find the set of spikes on each trial that contribute to this helix being detected (Fig. 4e, blue spikes), in the same manner as before, in simulated data (see for example Fig. 3, red spikes and red helix). Across extension trials when helix 2 was detected with high magnitude, the helix also occurred with similar timing (Fig. 4f). In other words, on extension trials when $||\mu_{2}||$ was large, $\text{Arg}(\mu_{2})$ was similar. This repeated structure can be visualized by overlaying the spikes across extension trials that align with helix 2 (Fig. 4f,g). When these spikes are overlaid across all extension perturbation trials with low transparency (Fig. 4g), dark regions indicate spikes that repeated reliably across trials, with some variability in timing (indicated by “smearing” in horizontal direction). On the other hand, when the spikes contributing to helix 2 across flexion trials are overlaid, no such structure is visible (Fig. 4h). On these trials, $||\mu_{2}||$ is small, and $\text{Arg}(\mu_{2})$ varies greatly; however, a different helix was reliably detected. Helix 8 was detected on flexion trials in this session, and occurred with reliable timing across trials. As before, this helix is readily apparent in the spikes across flexion trials (Fig. 4j), but is not visible in the spikes of extension trials (Fig. 4i). These results demonstrate how the ms-DHT can be used to detect robustly repeating and behaviourally relevant substructure in spike patterns, where the magnitude and phase of a helix contribution represents the amount and timing of such structure, respectively (as depicted in Fig.  3b).
[…]
I.4 Spike patterns in macaque prefrontal cortex
The recordings from two Utah arrays yield between 200-300 well-isolated single units in each session. In previous work, we found that sequences of single neuron activations encoded working memory content in this task (Fig. 5b) [39]. These sequences were composed of brief 100-200 ms elevations of firing rate well above baseline, and the position of each cell in the sequence was determined by the time it reached its maximum firing rate. The sequences had similar ordering and timing on trials of the same condition, and were separable between conditions. While we previously restricted to consider only one spiking event per cell on each trial, the ms-DHT now allows us to consider the complex pattern of all such elevations across all cells within a trial (Fig. 5c, Methods Sec.  II.10). This allows us to determine (1) whether we can replicate our previous results with this method and (2) whether including additional spiking events provides any additional insight.
@@ I.4 Spike patterns in macaque prefrontal cortex
To what do these “misplaced” trials correspond? When studying the navigation paths subjects took on these “misplaced” trials, the navigation paths are at first more similar to the trials with which the spike pattern clustered. Taking for example an individual trial with a target in the left column (encoded in blue), the spike pattern during navigation was more similar to trials with a target in the right column (Fig. 5g, blue dot in yellow cluster, circled in black). On that trial, the subject initially navigated to the right (Fig. 5h, black path compared to blue and yellow paths), reached the end of the arena without receiving a reward, then turned around and headed to the left to correct the mistake, eventually reaching the correct target (bold circle). These results extend to incorrect trials, where the similarity between these incorrect trials and the spike patterns on correct trials could explain the specific type of error made on individual trials (Fig. 5i,j). Finally, these results were consistent across trials, where we found that paths from misclassified trials were significantly more similar to paths in the predicted cluster than to paths in their true cluster (STATS) (see Supplement Fig. S3). These results show that spike patterns occurring during navigation could more closely reflect the navigation strategy than the end of the trajectory at the correct target (Fig. 5h,j, bold circle), and that analysis of spike patterns by the ms-DHT can predict the specific errors made on individual trials (see Methods Sec. II.10 for details).
@@ Discussion
In this work, we have introduced a novel framework for characterizing patterns of spiking activity across large populations of neurons. This approach represents the first complete mathematical language for spike patterns, which provides a simple way to describe, compare, and decompose patterns of spikes on a variety of timescales. We provide example applications to spiking data from macaque motor and prefrontal cortex recorded using Neuropixel probes and Utah arrays. These applications demonstrate how the multi-sample Discrete Helix Transform (ms-DHT) can be used to find behaviourally relevant structure in spike patterns that can then be directly related to the spikes on a single trial level.
@@ II.6 Dimensionality reduction procedure
In more detail, to compare $N_{T}$ spike patterns across $N$ cells, we begin by computing $\Psi_{H}(s_{i})$ for the spike pattern $s$ on each trial $i\in[1,N_{T}]$. We next compute the distance between each pair of spike patterns to produce the distance matrix $D\in\mathcal{M}_{N_{T}\times N_{T}}$ with $d_{i,j}:=||\Psi_{H}(s_{i})-\Psi_{H}(s_{j})||$. We then compute the eigenvalues $\lambda_{1},...\lambda_{N_{T}}$ and corresponding eigenvectors $\vec{v}_{1},....\vec{v}_{N_{T}}$, with eigenvalues ordered such that $|\lambda_{1}|\geq|\lambda_{2}|\geq...\geq\lambda_{N_{T}}$. Finally, we obtain the the $n$-dimensional projection $P:=D[\vec{v}_{1},...\vec{v}_{n}]$. This projection underlies the decoding in Fig. 4g,h and Fig. 5e, as well as the clustering in Fig. 5d,g,i.
@@ II.8 Significantly different helices
To determine whether helix contributions differed across trial conditions (as in Fig. 5n), we first computed a set of $1000$ bootstrapped centroids for each condition. The helix was considered to differ between conditions if these sets of centroids were non-overlapping (see for example Supplementary Figure 2b).
II.9 Application to Neuropixel data
Experimental details can be found at [45]. All procedures described below were approved by the Institutional Animal Care and Use Committee at Western University (Protocol 2022-028).
[…]
High-density Neuropixels probe recordings were performed (1.0 - 1 cm, 1.0 NHP - 1 cm, and 1.0 NHP - 4.5 cm). Neural recording targets were identified by registering the CT to a pre-surgery MRI, and identifying the 3D location of each brain area by warping segmentations from a composite macaque atlas to the individual MRI of each animal (NMT v2 [47, 48]). Neural data were recorded from Neuropixels probes using SpikeGLX. Neural data were processed using a custom processing pipeline [49]. For Monkey M, AP stream data was first drift corrected using spike localization and decentralized registration [50, 51] implemented in spikeinterface [52], which was able to accurately track vertical probe drift and correct it. Due to the large craniotomy, some of these recordings had large drift (0-250 um). Neural data were then processed with Kilosort 2.0 [53] to further stabilize recordings during spike sorting. For Monkey P, drift was minimal due to small craniotomies (drift 0-15 um), so the data was immediately processed using Kilosort 4.0 [53], including built-in drift correction. Single neurons were considered successfully recorded if they were flagged by Kilosort as single neurons using default parameters, and if they were stably recorded for the duration of the recording.
Data used in this application consisted of $100$ms of spiking activity immediately following the perturbation on each trial. Helices 2 and 8 were selected for visualization because the amplitude of their contributions reliably exceeded chance-level on only one type of perturbation. While a number of other helices also exhibited such selectivity (Supplemental Figure 2), selecting helices with lower spatial frequency aids visualization.
Decoding: The perturbation on single trials was decoded from the contributions of 2 and 10 of the most reliably occuring, perturbation-selective helices using a support vector machine (SVM). The SVM was trained on vectors of the real and imaginary parts (concatenated) of these helix contributions. Accuracy was reported with 10-fold cross-validation.
II.10 Application to Utah array data
@@
Experimental details and information about neuronal activation sequences can be found in [39]. Animal care and handling (i.e., basic care, animal training, surgical procedures, and experimental injections) were pre-approved by the University of Western Ontario Animal Care Committee. This approval ensures that federal (Canadian Council on Animal Care), provincial (Ontario Animals Research Act), regulatory bodies (e.g. CIHR/NSERC), and other national standards (CALAM) for the ethical use of animals are followed.
To extend these sequences to more complicated spike patterns, spike times were defined to be the center of a spiking event when a neuron exceeded $80\%$ of its peak firing rate. Epochs were divided by task timing: ITI = $1$ second before the appearance of the cue; “cue” = $3$ seconds following cue onset, while the visual cue is present on screen; “delay” = $2$ second memory delay; “response” = first $2$ seconds of navigation.
@@
We hypothesized that the paths taken by the animal on incorrectly clustered trials would more closely resemble the paths in the predicted cluster as opposed to the true cluster (i.e., if a blue trial is grouped into the yellow cluster, we expect the path on that trial to look more like a yellow path “predicted cluster” than a blue path “true cluster” - see Supplement Figs. S3a,b). We used the Fréchet distance to measure the distance between paths. As a baseline, we computed the distance between paths on trials correctly assigned to the same true cluster (see Supplement Fig. S3c “correct to true cluster”). This provides a value for the distance between the paths on all correct yellow column trials that fall into the yellow column cluster, for example. For each incorrectly clustered trial, we then computed the mean similarity to other paths in its predicted cluster, compared to other paths in its true cluster. As expected, paths on incorrectly clustered trials were significantly more similar to the paths on trials in the predicted cluster than to other trials of the same condition (true cluster) (p$<$0.003, see Supplement Fig. S3).
Together, this shows that transient spike patterns in the navigation epoch of the task closely reflect the navigation strategy taken by the subject on individual trials – so closely, in fact, that the observed spike pattern can predict the specific type of error made on individual trials.
II.11 Mathematical Details
II.11.1 Definitions
@@ II.11.2 Properties of $\Psi$
Data and Code Availability
Data and code used in this study will be made available at https://github.com/mullerlab/buschEAmathematical upon publication.